A single loop of wire of radius 3 cm carries a current of 2.8 A. What is the magnitude of B on the axis of the loop at the following positions?

1)the center of the loop

T

2)1 cm from the center

T

3)2 cm from the center

T

4)32 cm from the center

T
Answer #1

formual

B = uo i/ 2x

= 4 pi * 10^-7 i/ 2 r

(1)

r=0, B =4 pi * 10^-7 i/ 2 R

B = 4 pi * 10^-7 (2.8)/ 2 (0.03 ) = 5.86 * 10^-5 T

(2)

x = 1cm

B =4 pi * 10^-7 i R^2/ 2 ( R^2 + x^2 )^3/2

B = 4 pi * 10^-7 (2.8) ( 0.03)^2/ 2 (0.01^2 + 0.03^2 )^3/2 = 5.00 * 10^-5 T

(3)

x = 2cm

B =4 pi * 10^-7 i R^2/ 2 ( R^2 + x^2 )^3/2

B = 4 pi * 10^-7 (2.8) ( 0.03)^2/ 2 (0.02^2 + 0.03^2 )^3/2 = 3.37 * 10^-5 T

(4)

B =4 pi * 10^-7 i R^2/ 2 ( R^2 + x^2 )^3/2

B = 4 pi * 10^-7 (2.8) ( 0.03)^2/ 2 (0.32^2 + 0.03^2 )^3/2 = 4.76* 10^-8 T